A fixed volume of gas is initially at 180 degree C
Answer #1

PV = nRT

at fixed volume P1/T1 = P2/T2

=> T2 = P2*T1/P1 = 0.75Po*(180+273)/Po = 339.75 K = 67 degree celsius